# Topology-Aware Conformal Prediction for Stream Networks

### v1

@@ 1 Introduction
Stream networks represent a distinctive class of spatiotemporal graphs where data observations follow directional pathways and evolve dynamically over both space and time [15]. These networks are prevalent in various domains such as hydrology, transportation, and environmental monitoring, where data exhibit strong flow constraints [6, 12, 26, 45]. For example, in hydrology, river networks dictate the movement of water flow and pollutant dispersion [25], while in transportation, road and rail networks determine congestion and travel times [45, 44]. Understanding and modeling these networks are crucial for infrastructure planning, disaster response, and ecological conservation.
@@
A fundamental challenge in stream network analysis is predicting future observations and quantifying their uncertainty across multiple interconnected locations governed by network topology. Given the dynamic nature of these systems, accurate and reliable uncertainty quantification (UQ) is essential for risk assessment, decision-making, and resource allocation. For example, in transportation, estimating uncertainty in traffic volume forecasts across critical junctions enables optimal routing and congestion management [45]. However, the hierarchical dependencies, directional flow constraints, and evolving conditions inherent in stream networks introduce significant complexities in both predictive modeling and UQ.
@@
Recent advances in machine learning and statistical modeling have enhanced predictive accuracy for spatiotemporal data and enabled effective UQ with statistical guarantees [38, 9, 43]. In particular, conformal prediction (CP) has emerged as a powerful UQ framework, providing finite-sample validity guarantees under mild assumptions [27]. By constructing prediction sets with valid coverage probabilities, CP ensures that future observations fall within specified confidence intervals, enhancing reliability in decision-support systems [20, 17, 3, 42].
@@
Despite its success in various domains, traditional CP methods face significant limitations when applied to stream networks due to two key challenges: ($i$) Multivariate prediction: Unlike standard time-series predictions that focus on a single target variable, stream networks require joint predictions at multiple locations, where observations are highly interdependent. Applying CP independently at each location neglects network-wide dependencies, leading to inefficiencies in prediction set construction and potential loss of coverage guarantees. ($ii$) Intricate spatiotemporal flow constraints: Traditional CP assumes exchangeable data, an assumption that fails in stream networks due to directional flow constraints. While graph-based and spatial models account for topological relationships, stream networks exhibit unique dependency structures that neither conventional graph-based approaches nor purely data-driven models fully capture. Existing CP approaches either completely ignore dependencies without considering the spatiotemporal dynamics [28, 7] or attempt to learn dependencies solely from data without incorporating topological constraints [39, 30]. The former results in overly conservative or miscalibrated prediction sets, while the latter risks overfitting to specific network conditions, reducing generalizability.
@@
We provide a theoretical analysis of STACI, demonstrating that it maximizes prediction efficiency by reducing uncertainty set volume while maintaining valid coverage guarantees. To validate its effectiveness, we evaluate STACI on synthetic data with a stationary covariance matrix and real-world data with time-varying covariance, comparing its performance against state-of-the-art baseline methods [^1] Our codes are publicly available at https://github.com/fangxin-wang/STCP.. Both our theoretical and empirical results underscore the importance of the weighting parameter that balances data-driven insights with topology-induced knowledge, optimizing performance and enhancing predictive reliability in stream network applications.
@@ 2 Related Work
Stream networks, such as hydrology [15, 12], transportation networks [9], and environmental science networks [19], have been extensively studied due to their critical role in natural and engineered systems. Forecasting for stream network can be approached from two perspectives: as a graph prediction problem or as a multivariate time series prediction problem. In this work, we focus on the latter one, with the aim of predicting future data based on historical network data.
@@
Many approaches to stream network analysis relied on domain-specific statistical and physical models. Hoef et al. [12] introduced spatial stream network models for hydrology, emphasizing the importance of flow-connected relationships and spatial autocorrelation. The tail-up model [33] generalized spatial covariance structures to stream networks by weighting observations based on flow connectivity. Recent advances in machine learning have spurred innovative approaches for modeling stream networks, particularly through graph-based frameworks that leverage their inherent spatio-temporal (ST) graph dynamics. Within ST graph forecasting, many current approaches emphasize point estimation, which aims to predict the most likely future values [14, 8].
@@
While effective and widely adopted, models without uncertainty quantification often lack considerations for reliability, posing limitations particularly in safety-critical scenarios. To address this, some studies [38, 46, 29, 37, 26] turn to explore interval prediction, which ensures that prediction intervals cover the ground-truth values with a pre-defined high probability, offering a more reliable alternative. Among these approaches, the majority of studies employ Bayesian methods to construct prediction intervals for ST forecasting problems [36]. These methods commonly utilize Monte Carlo Dropout [38, 26] or Probabilistic Graph Neural Networks [46, 37]. However, the performance of Bayesian methods has been found to be sensitive to the choice of prediction models and priors, particularly the type of probabilistic distributions [37]. To address these limitations, classic Frequentist-based methods, such as quantile regression and conformal prediction, have been employed, which generally offer more robust coverage across data and model variations.
@@
Conformal prediction (CP) [35] has recently gained traction across multiple domains, including graphs [5, 13, 21] and multi-dimensional time series [30, 23, 40]. Since ST graphs can naturally be viewed as a special case of multi-dimensional time series, we focus on CP methods designed for this setting. Sun and Yu [30] assumes that data samples for each entire time series are drawn independently from the same distribution, while Messoudi et al. [23] assumes exchangeability in the data. Both approaches fail to capture the complex temporal and spatial dependencies inherent in ST graphs, limiting their applicability. Xu et al. [40] construct ellipsoidal prediction regions for non-exchangeable multi-dimension time series, but their model neglects the inherent graph structure embedded within the multi-dimensional time series and overlook scenarios where the error process (see Equation 3.1) is non-stationary, a prevalent feature in real-world data. Section B provides a taxonomy of existing CP methods, highlighting our unique positioning within the CP literature. To the best of our knowledge, no previous work has specifically tailored CP for stream networks or other spatio-temporal graphs, reinforcing the novelty and importance of our contribution to this domain.
3 Problem Setup and Preliminaries
@@
Consider a stream network ${\mathcal{G}}$ with fixed flow direction at time $t\in\{1,\dots,T\}$, with $I$ observational sites indexed by =$\mathcal{I}=\{1,\dots,I\}$. Let $\mathcal{L}\subset\mathbb{R}^{2}$ denote the set of all geolocations on the network, and let the geographical location of site $i\in\mathcal{I}$ be represented as $\ell_{i}\in\mathcal{L}$. The stream network consists of segments $\{r_{j}\subset\mathcal{L},j\in\mathcal{J}\}$, where $\mathcal{J}$ is the index set of all stream segments. Each site $i\in\mathcal{I}$ is located within a specific segment $r_{j}$ for some $j\in\mathcal{J}$, and a segment may contain multiple or no observational sites. For any location $u\in\mathcal{L}$, we define $\wedge u$ as the set of all upstream segments of location $u$, and $\vee u$ as the set of all downstream segments of location $u$. The hydrologic distance between two locations $v,u\in\mathcal{L}$, denoted as $d(v,u)$, is the distance measured along the stream. If $v$ and $u$ belong to the same segment $s_{i}$, $d(v,u)$ is simply the Euclidean distance between $v$ and $u$. See Figure 4 for an illustration.
@@
This objective can be achieved using split conformal prediction (CP) [35], a widely used statistical framework for uncertainty quantification. Split CP operates by first partitioning the data into a training set and a calibration set. The prediction model $\hat{f}$ is trained exclusively on the training set. To assess the reliability of predictions, a nonconformity score is computed, which quantifies the deviation of each calibration sample from the ground truth. Given a target confidence level $\alpha$, the method determines the $(1-\alpha)$-th quantile of the nonconformity scores from the calibration data. This quantile is then used to adjust $\hat{f}$’s predictions for test samples, ensuring the constructed prediction sets maintain valid coverage. Under the assumption that the calibration and test data are exchangeable, the resulting prediction sets are guaranteed to achieve a coverage rate of at least $1-\alpha$ on the test data.
@@ 4 Proposed Framework
4.1 Topology-aware Nonconformity Score
@@
To account for the intricate spatio-temporal dependencies, we consider a general class of nonconformity score functions based on the Mahalanobis distance [16]:
$s(X_{t},Y_{t})\coloneqq\hat{\epsilon}_{t}^{\top}A\hat{\epsilon}_{t},\quad\forall t\in\mathcal{D}_{\text{cal}},$ (4.1)
@@
$A\coloneqq(1-\lambda)\hat{\Sigma}^{-1}_{n}+\lambda\hat{\Sigma}^{-1}_{{\mathcal{G}}}.$ (4.2)
@@
Unlike the method proposed by [39], which relies solely on the sample covariance estimate, this formulation incorporates the underlying topology of the stream network. By balancing data-driven and structural information, it provides a more robust covariance estimation, leading to better prediction efficiency without sacrificing coverage validity.
@@
We develop a novel method to estimate the topology-induced covariance $\hat{\Sigma}_{{\mathcal{G}}}$ used in Equation 4.2 by assuming the observations on the stream network can be captured by a tail-up model [12, 33, 11]. The tail-up model is formally defined as follows:
Definition 4.1 (Tail-up model).
@@
[…]
where $\wedge u$ denotes all the segments that are the upstream of $u$. Here, $\mu(u)$ is the deterministic mean at $u$, and $m(v-u)$ is a moving average function capturing the influence from upstream location $v$ to $u$. Both $w(v)$ and $w(u)$ are weights that satisfy the additivity constraint such that the variance remains constant across sites.
We note that the tail-up model only requires the assumptions of ergodicity and spatial stationarity [32], which is highly flexible and can be broadly applied to a wide range of stream network data. Also, the choice of the moving average function $m(\Delta)$ remains adaptable, as long as it has a finite volume, allowing the model to accommodate different spatial structures effectively.
To estimate $\hat{\Sigma}_{{\mathcal{G}}}$, we model $B(v)$ using Brownian motion and adopt an exponential moving average function for $m(\Delta)=\beta\exp(-\Delta/\phi)$. Therefore, the topology-induced covariance between any two locations $u,v$ can be expressed as follows (See the proof in Lemma A.5):
$\hat{\Sigma}_{{\mathcal{G}}}(u,v)=\begin{cases}\sigma^{2}\sqrt{\frac{w(u)}{w(v)}}\exp\left(-\dfrac{d(u,v)}{\phi}\right)&\text{if }u\to v,\\ 0&\text{otherwise},\end{cases}$ (4.4)
@@
Intuitively, the resulting covariance structure captures how information propagates along the stream network. The exponential decay function in Equation 4.4 ensures that the influence of an upstream location $u$ on a downstream location $v$ diminishes as their hydrologic distance $d(u,v)$ increases. The weighting term $\sqrt{w(u)/w(v)}$ further adjusts this influence based on flow intensity, reflecting the fact that locations with stronger flow connections exert a greater impact on each other. This formulation naturally aligns with real-world stream dynamics, where observations at one site are more strongly correlated with those from nearby upstream sources, while distant or disconnected locations exhibit little to no dependence.
4.2 Adaptive Uncertainty Set Construction
We construct a spatio-temporally adaptive prediction set for a new observed history $X_{T+1}$ using our proposed nonconformity score, defined in Equation 4.1, as follows:
@@ 4.2 Adaptive Uncertainty Set Construction
where $\hat{Q}_{1-\alpha}$ is the $(1-\alpha)$-th quantile of the empirical cumulative distribution function of $\{s(X_{t},Y_{t}),t\in\mathcal{D}_{cal}\}$. The complete STACI algorithm is outlined in Algorithm 1.
Algorithm 1 Offline STACI
Output: Prediction set $\mathcal{C}(X_{T+1};\alpha)$.
1:   // Training
2:  $\hat{f}\leftarrow$ Fit $f$ using $\mathcal{D}\setminus\mathcal{D}_{\text{cal}}$;
3:   // Calibration
4:  $\mathcal{E}\leftarrow\{\hat{\epsilon}_{t}=Y_{t}-\hat{f}(X_{t})-\frac{\sum_{t\in\mathcal{D}_{\text{cal}}}(Y_{t}-\hat{f}(X_{t}))}{|\mathcal{D}_{\text{cal}}|}\}_{t\in\mathcal{D}_{\text{cal}}}$;
5:  $\hat{\Sigma}_{n}\leftarrow\sum_{t\in\mathcal{D}_{\text{cal}}}\hat{\epsilon}_{t}\hat{\epsilon}_{t}^{\top}/(n-1)$ given $\mathcal{E}$;
6:  $\hat{\Sigma}_{\mathcal{G}}\leftarrow$ Compute (4.4) for $(\ell_{i},\ell_{i^{\prime}}),\forall i,i^{\prime}\in\mathcal{I}$ given ${\mathcal{G}}$;
7:  $A\leftarrow\lambda\hat{\Sigma}_{\mathcal{G}}^{-1}+(1-\lambda)\hat{\Sigma}_{n}^{-1}$;
8:  $\mathcal{S}\leftarrow\{\hat{\epsilon}_{t}^{\top}A\hat{\epsilon}_{t}\}_{t\in\mathcal{D}_{\text{cal}}}$ given $\mathcal{E}$;
9:  $\hat{Q}_{1-\alpha}\leftarrow$ Compute $\frac{\lceil(1-\alpha)(n+1)\rceil}{n}$-th quantile given $\mathcal{S}$;
10:   // Testing
11:  $\mathcal{C}(X_{T+1};\alpha)\leftarrow\{y:s(X_{T+1},y)\leq\hat{Q}_{1-\alpha}\}$;
To account for potential temporal distribution shifts in the predictive error of Equation 3.1, we adopt the Adaptive Conformal Inference (ACI) framework proposed in [10]. This approach dynamically updates the confidence level $\alpha_{t}$ over time, ensuring that the prediction set remains responsive to evolving data distributions. Specifically, we iteratively update $\alpha_{t}$, and reconstruct the prediction set $\mathcal{C}(X_{t},\alpha_{t})$ accordingly. At the initial test time $T+1$, the confidence level is set as $\alpha_{T+1}=\alpha$. For subsequent time steps $t>T+1$, $\alpha_{t}$, we update $\alpha_{t}$ with a step size $\gamma\geq 0$ as follows:
$\alpha_{t+1}=\alpha_{t}+\gamma(\alpha-\mathds{1}\{Y_{t}\notin\mathcal{C}(X_{t};\alpha_{t})\}),\quad\forall t\geq T+1.$ (4.5)
The rationale behind ACI is that if the prediction set fails to cover the true value at time $t$, the effective error level is reduced, leading to a wider prediction interval at time $t+1$, thereby increasing the likelihood of coverage. A larger step size $\gamma$ makes the method more responsive to observed distribution shifts but also introduces greater fluctuations in $\alpha_{t}$. When $\gamma=0$, the method reduces to standard (non-adaptive) conformal prediction.

### v2

@@ 1 Introduction
Stream networks represent a distinctive class of spatiotemporal graphs where data observations follow directional pathways and evolve dynamically over both space and time isaak2014applications. These networks are prevalent in various domains such as hydrology, transportation, and environmental monitoring, where data exhibit strong flow constraints cressie2006spatial; hoef2006spatial; qian2023uncertainty; zhu2021spatio; sheth2023streams. For example, in hydrology, river networks dictate the movement of water flow and pollutant dispersion qadir2008spatio, while in transportation, road and rail networks determine congestion and travel times zhu2021spatio; zhou2024hierarchical. Understanding and modeling these networks are crucial for infrastructure planning, disaster response, and ecological conservation.
@@
A fundamental challenge in stream network analysis is predicting future observations and quantifying their uncertainty across multiple interconnected locations governed by network topology. Given the dynamic nature of these systems, accurate and reliable uncertainty quantification (UQ) is essential for risk assessment, decision-making, and resource allocation. For example, in transportation, estimating uncertainty in traffic volume forecasts across critical junctions enables optimal routing and congestion management zhu2021spatio. However, the hierarchical dependencies, directional flow constraints, and evolving conditions inherent in stream networks introduce significant complexities.
@@
Recent advances in machine learning and statistical modeling have enhanced predictive accuracy for spatio-temporal data and enabled effective UQ with statistical guarantees wu2021trafficquantifying; gao2023spatiotemporal; zheng2024optimizing. In particular, conformal prediction (CP) has emerged as a powerful UQ framework, providing finite-sample validity guarantees under mild assumptions shafer2008tutorial. By constructing prediction sets with valid coverage probabilities, CP ensures that future observations fall within specified confidence intervals, enhancing reliability in decision-support systems lei2018distribution; kivaranovic2020conformal; cauchois2021knowing; zargarbashi2023conformal.
@@
Despite its success in various domains, traditional CP methods face significant limitations when applied to stream networks due to two key challenges: ($i$) Multivariate prediction: Unlike standard time-series predictions that focus on a single target variable, stream networks require joint predictions at multiple locations, where observations are highly interdependent. Applying CP independently at each location neglects network-wide dependencies, leading to inefficiencies in prediction set construction and potential loss of coverage guarantees. ($ii$) Intricate spatio-temporal flow constraints: Traditional CP assumes exchangeable data, an assumption that fails in stream networks due to directional flow constraints. While graph-based and spatial models account for topological relationships, stream networks exhibit unique dependency structures that neither conventional graph-based approaches nor purely data-driven models fully capture. Existing CP approaches either completely ignore dependencies without considering the spatio-temporal dynamics stankeviciute2021conformal; diquigiovanni2021distribution or attempt to learn dependencies solely from data without incorporating topological constraints xu2023conformal; sun2024copula. The former results in overly conservative or miscalibrated prediction sets, while the latter risks overfitting to specific network conditions, reducing generalizability.
@@
We provide a theoretical analysis of STACI, demonstrating that it maximizes prediction efficiency by reducing uncertainty set volume while maintaining valid coverage guarantees. To validate its effectiveness, we evaluate STACI on synthetic data with a stationary covariance matrix and real-world data with time-varying covariance, comparing its performance against state-of-the-art baseline.  [^1] Our code is publicly available at https://github.com/fangxin-wang/STCP. Both our theoretical and empirical results underscore the importance of the weighting parameter that balances data-driven insights with topology-induced knowledge, optimizing performance and enhancing predictive reliability in stream network applications.
@@ 2 Related Works
Stream networks, such as hydrology isaak2014applications; hoef2006spatial; sheth2023streams, transportation networks gao2023spatiotemporal; liu2023largest, and environmental science networks launay2015calibrating, have been extensively studied due to their critical role in natural and engineered systems. Forecasting for stream network can be approached from two perspectives: as a graph prediction problem or as a multivariate time series prediction problem. In this work, we focus on the latter one, with the aim of predicting future data based on historical network data.
@@
Many approaches to stream network analysis rely on domain-specific statistical and physical models. hoef2006spatial introduced spatial stream network models for hydrology, emphasizing the importance of flow-connected relationships and spatial autocorrelation. The tail-up model ver2010moving generalized spatial covariance structures to stream networks by weighting observations based on flow connectivity. Recent advances in machine learning have spurred innovative approaches for modeling stream networks for point forecasts, particularly through graph-based frameworks that leverage their inherent spatio-temporal (ST) graph dynamics huang2014deep; du2020hybrid.
@@
While effective and widely adopted, models without uncertainty quantification often lack considerations for reliability, posing limitations particularly in safety-critical scenarios. To address this, some studies wu2021trafficquantifying; zhuang2022trafficuncertainty; sun2022conformal; wang2023trafficuncertainty; qian2023uncertainty turn to explore interval prediction, which ensures that prediction intervals cover the ground-truth values with a pre-defined high probability, offering a more reliable alternative. Among these approaches, the majority of studies employ Bayesian methods to construct prediction intervals for ST forecasting problems wang2024uncertainty. These methods commonly utilize Monte Carlo Dropout wu2021trafficquantifying; qian2023uncertainty or Probabilistic Graph Neural Networks zhuang2022trafficuncertainty; wang2023trafficuncertainty. However, the performance of Bayesian methods has been found to be sensitive to the choice of prediction models and priors, particularly the type of probabilistic distributions wang2023trafficuncertainty. To address these limitations, classic Frequentist-based methods have been employed, such as conformal prediction, which generally offer more robust coverage across data and model variations.
@@
Conformal prediction (CP) vovk2005algorithmic has recently gained significant attention across various domains, including graph-structured data clarkson2023distribution; huang2023uncertainty; lunde2023conformal and multi-dimensional time series sun2024copula; messoudi2021copula; xuconformal. Existing CP methods for graphs clarkson2023distribution; huang2023uncertainty; lunde2023conformal have primarily addressed node/edge classification and ranking tasks, where UQ concerns discrete labels or scores on static graph structures. In contrast, our work focuses on a fundamentally different problem: UQ for spatio-temporal forecasting on dynamic networks. Given that spatio-temporal graphs can be naturally formulated as a special case of multivariate time series, we discuss related CP methodologies developed for the latter setting sun2024copula; messoudi2021copula; xuconformal. sun2024copula assumes that data samples for each entire time series are drawn independently from the same distribution, while messoudi2021copula assumes exchangeability in the data. Both approaches fail to capture the complex temporal and spatial dependencies inherent in ST graphs, limiting their applicability. xuconformal construct ellipsoidal prediction regions for non-exchangeable multi-dimensional time series, but their model neglects the inherent graph structure embedded within the multi-dimensional time series and overlook scenarios where the error process (see eq.˜1) is non-stationary, a prevalent feature in real-world data. Section B in the Appendix provides a taxonomy of existing CP methods, highlighting our unique positioning within the CP literature. To the best of our knowledge, no previous work has specifically tailored CP for stream networks or other spatio-temporal graphs, reinforcing the novelty to this domain.
3 Problem
@@
Consider a stream network $\mathcal{G}$} with fixed flow direction at time $t\in\{1,\dots,T\}$, with $I$ observational sites indexed by $\mathcal{I}=\{1,\dots,I\}$. Let $\mathcal{L}\subset\mathbb{R}^{2}$ denote the set of all geolocations on the network, and let the geographical location of site $i\in\mathcal{I}$ be represented as $\ell_{i}\in\mathcal{L}$. The stream network consists of segments $\{r_{j}\subset\mathcal{L},j\in\mathcal{J}\}$, where $\mathcal{J}$ is the index set of all stream segments. Each site $i\in\mathcal{I}$ is located within a specific segment $r_{j}$ for some $j\in\mathcal{J}$, and a segment may contain multiple or no observational sites. For any location $u\in\mathcal{L}$, we define $\wedge u$ as the set of all upstream segments of location $u$, and $\vee u$ as the set of all downstream segments of location $u$. The hydrologic distance between two locations $v,u\in\mathcal{L}$, denoted as $d(v,u)$, is the distance measured along the stream. If $v$ and $u$ belong to the same segment $r_{i}$, $d(v,u)$ is simply the Euclidean distance between $v$ and $u$. See Figure 1 for an illustration.
@@
This objective can be achieved using split conformal prediction (CP) vovk2005algorithmic, a widely used framework for uncertainty quantification. Split CP operates by first partitioning the data into a training set and a calibration set. The prediction model $\hat{f}$ is trained exclusively on the training set. To assess the reliability of predictions, a nonconformity score is computed, which quantifies the deviation of each calibration sample from the ground truth. Given a target confidence level $\alpha$, the method determines the $(1-\alpha)$-th quantile of the nonconformity scores from the calibration data. This quantile is then used to adjust $\hat{f}$’s predictions for test samples, ensuring the constructed prediction sets maintain valid coverage. Under the assumption that the calibration and test data are exchangeable, the prediction sets are guaranteed to achieve a coverage rate of at least $1-\alpha$ on the test data.
@@ 4 Proposed Framework
Topology-aware Nonconformity Score
@@
To account for the intricate spatio-temporal dependencies, we consider a general class of nonconformity score functions based on the Mahalanobis distance pmlr-v230-katsios24a:
$s(X_{t},Y_{t})\coloneqq\hat{\epsilon}_{t}^{\top}A\hat{\epsilon}_{t},\quad\forall t\in\mathcal{D}_{\text{cal}},$ (2)
@@
$A\coloneqq(1-\lambda)\hat{\Sigma}^{-1}_{n}+\lambda\hat{\Sigma}^{-1}_{{\mathcal{G}}}.$ (3)
@@
Unlike the method proposed by xu2023conformal, which relies solely on the sample covariance estimate, this formulation incorporates the underlying topology of the stream network. By balancing data-driven and structural information, it provides a more robust covariance estimation, leading to better prediction efficiency without sacrificing coverage validity.
@@
We develop a novel method to estimate the topology-induced covariance $\hat{\Sigma}_{{\mathcal{G}}}$ used in eq.˜3 by assuming the observations on the stream network can be captured by a tail-up model hoef2006spatial; ver2010moving; higdon2022non. The tail-up model is formally defined as follows:
Definition 1 (Tail-up model).
@@
$Y(u)=\mu(u)+\int_{\wedge u}m(v-u)\sqrt{\frac{w(v)}{w(u)}}dB(v),$ (4)
where $\wedge u$ denotes all the segments that are the upstream of $u$. Here, $\mu(u)$ is the deterministic mean at $u$, and m(vu)m(v-u) is a moving average function capturing the influence from upstream location $v$ to $u$. Both $w(v)$ and $w(u)$ are weights that satisfy the additivity constraint such that the variance remains constant across sites.
We note that the tail-up model only requires the assumptions of ergodicity and spatial stationarity ver1993multivariable, which is highly flexible and can be broadly applied to a wide range of stream network data. Also, the choice of the moving average function $m(\Delta)$ remains adaptable, as long as it has a finite volume, allowing the model to accommodate different spatial structures effectively.
To estimate $\hat{\Sigma}_{{\mathcal{G}}}$, we model $B(v)$ using Brownian motion and adopt an exponential moving average function for $m(\Delta)=\beta\exp(-\Delta/\phi)$. Therefore, the topology-induced covariance between any two locations $u,v$ can be expressed as follows (See Lemma  5 in Section A of the Appendix):
$\hat{\Sigma}_{\mathcal{G}}(u,v)=\sigma^{2}\sqrt{\frac{w(u)}{w(v)}}\exp\bigl(-d(u,v)/\phi\bigr)\,\text{, \;if $u\to v$},$ (5)
@@
Intuitively, the covariance structure reflects how information propagates along the stream network. The exponential decay in eq.˜5 models diminishing influence with increasing hydrologic distance $d(u,v)$, while the weight term $\sqrt{w(u)/w(v)}$ modulates this effect based on flow intensity. This formulation naturally aligns with real-world stream dynamics, where proximal upstream sites exert stronger influence than distant or disconnected ones.
Adaptive Uncertainty Set Construction
We construct a spatio-temporally adaptive prediction set for a new observed history $X_{T+1}$ using our proposed nonconformity score, defined in eq.˜2, as follows:
@@
where $\hat{Q}_{1-\alpha}$ is the $(1-\alpha)$-th quantile of the empirical cumulative distribution function of $\{s(X_{t},Y_{t}),t\in\mathcal{D}_{cal}\}$.
To account for potential temporal distribution shifts in the predictive error of eq.˜1, we adopt the Adaptive Conformal Inference (ACI) framework proposed in gibbs2021adaptive. This approach dynamically updates the confidence level $\alpha_{t}$ over time, ensuring that the prediction set remains responsive to evolving data distributions. Specifically, we iteratively update $\alpha_{t}$, and reconstruct the prediction set $\mathcal{C}(X_{t},\alpha_{t})$ accordingly. At the initial test time $T+1$, the confidence level is set as $\alpha_{T+1}=\alpha$. For subsequent time steps $t>T+1$, $\alpha_{t}$, we update $\alpha_{t}$ with a step size $\gamma\geq 0$ as follows:
$\alpha_{t+1}=\alpha_{t}+\gamma(\alpha-\mathds{1}\{Y_{t}\notin\mathcal{C}(X_{t};\alpha_{t})\}),\quad\forall t\geq T+1.$ (6)
The rationale behind ACI is that if the prediction set fails to cover the true value at time $t$, the effective error level is reduced, leading to a wider prediction interval at time $t+1$, thereby increasing the likelihood of coverage. A larger step size $\gamma$ makes the method more responsive to observed distribution shifts but also introduces greater fluctuations in $\alpha_{t}$. When $\gamma=0$, the method reduces to standard conformal prediction with fixed $\alpha$. The detailed analysis of coverage guarantee of Adaptive Uncertainty Set construction is discussed in Section C in Appendix.